# OpenEarthMap-SAR: A Benchmark Synthetic Aperture Radar Dataset for Global High-Resolution Land Cover Mapping

### v1

@@ II-B Classes, Annotations, and Data Split
  • Real labels (1/5 per region) only: The model will be trained exclusively on the real labels (1 or 5 per folder).
For the semantic segmentation task, we will evaluate the performance with Optical, Optical, and Optical+SAR data combinations to assess the models across different modalities.
@@ II-C Comparison with other datasets
The OpenEarthMap-SAR dataset stands out in comparison to other high-resolution SAR benchmark datasets in terms of its region coverage and class diversity. While datasets like FU-SAR [16], GF-3 Building [17], and BDD [18] focus on specific tasks such as Land Cover Mapping (LC), Building Extraction (BE), and Change Detection (CD) with smaller class counts, OpenEarthMap-SAR offers a broader range of classes (8 in total) and covers more regions (35), spanning across 3 countries. It also presents a resolution range of 0.15–0.5 meters, placing it in the sub-meter level, similar to datasets such as SpaceNet 6 [19] and BRIGHT [3]. However, the OpenEarthMap-SAR dataset surpasses other datasets in terms of its extensive coverage, both geographically and in the number of regions, making it a valuable resource for studying land cover patterns in SAR imagery. Despite the smaller number of classes compared to certain other datasets like FU-SAR, which includes 4 classes, OpenEarthMap-SAR’s diverse geographic scope and multiple region coverage provide a unique advantage for comprehensive land cover analysis

### v2

@@ II-B Classes, Annotations, and Data Split
  • Real labels (1/5 per region) only: The model will be trained exclusively on the real labels (1 or 5 per region).
For the semantic segmentation task, we will evaluate the performance with SAR, Optical, and Optical+SAR data combinations to assess the models across different modalities.
@@ II-C Comparison with other datasets
The OpenEarthMap-SAR dataset stands out in comparison to other high-resolution SAR benchmark datasets in terms of its region coverage and class diversity. While datasets like FU-SAR [16], GF-3 Building [17], and BDD [18] focus on specific tasks such as Land Cover Mapping (LC), Building Extraction (BE), and Change Detection (CD) with smaller class counts, OpenEarthMap-SAR offers a broader range of classes (8 in total) and covers more regions (35), spanning across 3 countries. It also presents a resolution range of 0.15–0.5 meters, placing it in the sub-meter level, similar to datasets such as SpaceNet 6 [19] and BRIGHT [3]. However, the OpenEarthMap-SAR dataset surpasses other datasets in terms of its extensive coverage, both geographically and in the number of regions, making it a valuable resource for studying land cover patterns in SAR imagery. Despite the smaller number of classes compared to certain other datasets like FU-SAR, which includes 4 classes, OpenEarthMap-SAR’s diverse geographic scope and multiple region coverage provide a unique advantage for comprehensive land cover analysis.